Merry is a trucker and needs to keep track of her mileage for business purposes. She begins a trip with an odometer reading of r miles and ends the trip with an odometer reading of t miles.
a. Express the total of miles covered algebraically.
b. If the tip takes h hours of driving, express the average speed algebraically.
c. If the truck used g gallons of gas during the trip, express the mpg the truck averaged during the trip algebraically.

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Merry is a trucker and needs to keep track of her mileage for business purposes. She begins a trip with an odometer reading of r miles and ends the trip with an odometer reading of t miles.
a. Express the total of miles covered algebraically.
Ans: t-r miles
----------------------
b. If the tip takes h hours of driving, express the average speed algebraically.
Ans: (t-r)/h miles/hour
-----------------------
c. If the truck used g gallons of gas during the trip, express the mpg the truck averaged during the trip algebraically.
Ans: [(t-r)/g] miles/gallon
=============
Cheers,
Stan H.
-------------

Answer by ewatrrr(24785)

trip begins with an odometer reading of r miles and ends the trip with an odometer reading of t miles.
.....
total of miles covered algebraically: (t-r)
h hours of driving, express the average speed : (t-r)/h
used g gallons of gas during the trip, average mpg = (t-r)/g
.......
mph = mi/hr and mpg = mi/gal